UnknownDakota:

Probability Questions... A player gets to throw 4 darts at the target shown. Assuming the player will always hit the target, the probability of hitting an odd number three times is ______ times more than the probability of hitting an even number three times.

UnknownDakota:

1 attachment
Vocaloid:

there are three odd numbers (1,3,5) and two even numbers (2,4) out of 5 total possibilities the probability of hitting one odd number is therefore 3/5 to get three in a row, we just multiply this by itself three times, so P(three odd numbers in a row) = (3/5)(3/5)(3/5) same logic with getting three even numbers in a row P(three even numbers in a row) = (2/5)(2/5)(2/5)

Vocaloid:

so you would just figure out how many times more (3/5)^3 is compared to (2/5)^3

UnknownDakota:

2.25

UnknownDakota:

i have another question if youd like to help

Vocaloid:

hm I think i got something slightly different for (3/5)^3 / (2/5)^3

Vocaloid:

actually hmmm

Vocaloid:

since the person gets to throw four darts I think the calculation becomes a bit more complex

UnknownDakota:

it was 2.25 apparently on plato

Vocaloid:

ah I see

Vocaloid:

if you have another one I'll take a look

UnknownDakota:

even my teacher is having trouble

1 attachment
UnknownDakota:

this is the last question for me to graduate. unless i get it wrong

Vocaloid:

hmm I really don't feel confident about this solution, my best guess is <XOB = 2 times <QON, personally I'd try to get a second opinion :/ maybe I can ask someone

UnknownDakota:

please do

Vocaloid:

sure

UnknownDakota:

it was correct

UnknownDakota:

<QON = 55 so 55 * 2 = <XOB <XOB Is 110